A spherical conductor having radius 20cm is charged such that at a distance 5cm from the center v=20v find the potential difference from the center also find excess proton in the conductor

Grade:12

1 Answers

Eshan
5 years ago
Dear student,

All the charges on a spherical conductor lies on its surface and no electric exists inside. Hence potential at all points inside the sphere is equal to that on the surface of the conductor.

Hence potential at the given point=potential on surface of sphere=\dfrac{kQ}{R}=20
\implies Q=4.44\times 10^{-10}C
Hence excess of protons in the conductor=\dfrac{Q}{e}=2.77\times 10^9